# On the Algorithmic Bias of Aligning Large Language Models with RLHF: Preference Collapse and Matching Regularization

### v1

@@ 3 Preference Matching RLHF
We start by presenting a motivating example. Given a reward model, consider the simple scenario of optimizing the expected reward without any regularization. For a question $x$, formally called a prompt, let $y_{1}$ and $y_{2}$ be two possible responses. Imagine 51% of human labelers prefer $y_{1}$ over $y_{2}$ for this binary comparison. Assuming that the reward model accurately reflects this 51% versus 49% preference between the two responses, it will assign a slightly higher reward $r(x,y_{1})$ to the first response than $r(x,y_{2})$ to the second one. Maximizing the reward without any regularization would lead the LLM to exclusively prefer the majority opinion and completely disregard any minority opinions.
@@
Let $\phi^{\star}$ be an optimal solution to the unregularized reward maximization problem [^4] We assume sufficient expressivity of $\pi_{\phi}$ and that the expected reward is maximized exactly, as detailed in Hypotheses 3.1 and 3.2 in Section 3.1.
$\max_{\phi}\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}r(x,y),$
where the expectation is over the randomness of both $x$ and $y$ following the conditional distribution $\pi_{\phi}(\cdot|x)$. For a fixed $x$, with probability one, $\pi_{\phi^{\star}}(y|x)$ outputs some response $y^{\star}$ with the highest reward $r(x,y^{\star})$.
Although this result is almost trivial, [^5] Its proof is based on the fact that a weighted average is maximized if all weights are assigned to the highest values. it implies that regularization is necessary to match the preferences of the reward model. Formally, we consider adding a regularizer $R(\pi_{\phi}(y|x))$ to the reward maximization problem:
@@
Our goal in this section is to find $R(\cdot)$ by first obtaining necessary conditions ensuring that the global solution to (3.1) is algorithmically unbiased.
In Section 3.3, we further extend our attention to a regularizer $R$ that depends on both the prompt $x$ and response $y$. This extended scope particularly allows the discussion to cover the KL divergence used in the standard RLHF (2.3), in which case we can let
$R_{x,y}(\cdot)=-\beta\log\bigg{(}\frac{\cdot}{\pi_{\textnormal{ref}}(y|x)}\bigg{)}$
since $\pi_{\textnormal{ref}}(y|x)$ can be computed from $x$ and $y$.
Definition 3.2 (Preference Matching).
We say that $\pi(y|x)$ is a preference matching (PM) [^6] Note that preference matching is not abbreviated in Section 1 for illustration purposes, which is the same rule for BTL but not for LLM. policy with respect to a reward model $r(x,y)$ if the probability distribution of its output for any prompt $x$ matches $r(x,y)$ under the PL model. That is,
$\pi(y|x)=\frac{\mathrm{e}^{r(x,y)}}{\sum_{y^{\prime}}\mathrm{e}^{r(x,y^{\prime})}}.$ (3.2)
Remark 3.3.
From a practical perspective, we only require (3.2) to hold for prompts where there is no single correct answer and the response depends on personal opinions or preferences. Another reason for not requiring it to hold for all prompts is that the reward model might not generalize well for some prompts.
@@ 3.1 Preference Matching Differential Equation
Assuming that $R$ is sufficiently differentiable, we aim to find the necessary condition on $R$ that leads to PM regularization based on (3.5). Consider the Lagrangian of (3.5): [^7] Here, we omit the the constraints $\pi_{i}\geq 0$ since they will be satisfied automatically.
$L=\sum_{i=1}^{k}\pi_{i}(r_{i}+R(\pi_{i}))+\lambda\left(\sum_{i=1}^{k}\pi_{i}-1\right).$
The first-order conditions
$\frac{\partial L}{\partial\pi_{i}}=r_{i}+R(\pi_{i})+\pi_{i}R^{\prime}(\pi_{i})+\lambda=0$
must be satisfied by the PM policy $\pi^{\star}$, which obeys (3.2). Thus, we get
@@
for all $i$, where $\pi^{\star}_{i}:=\pi^{\star}(y_{i}|x)=\frac{\mathrm{e}^{r_{i}}}{\sum_{j=1}^{k}\mathrm{e}^{r_{j}}}$. Conversely, we can write
$r_{i}=\log\left(\pi^{\star}_{i}\sum_{j=1}^{k}\mathrm{e}^{r_{j}}\right)=\log\pi_{i}^{\star}+c,$
where $c=\log(\sum_{j=1}^{k}\mathrm{e}^{r_{j}})$ does not depend on $i=1,\ldots,k$. This expression of $r_{i}$ together with (3.6) gives
@@ 3.2 Preference Matching Regularization
where $C_{1}$ and $C_{2}$ are constants. In words, $\pi^{\star}(y|x)\equiv\mathrm{e}^{r(x,y)}/\sum_{y^{\prime}}\mathrm{e}^{r(x,y^{\prime})}$ is the global solution to (3.4) only if $R$ takes the form above.
It is important to recognize that the form of $R$ here does not depend on the specific probabilities $\pi_{i}^{\star}$’s, which implicitly depend on the prompt $x$. This is a convenient fact that facilitates the implementation of the PM regularization. If there were such dependence, it would be very complicated to compute $R$ since it would require knowledge of $\pi_{i}^{\star}$’s from the reward model.
Notably, the derivation of (3.8) is conditional on $x$, which implies that the constants $C_{1}$ and $C_{2}$ can depend on $x$. Taken together, our discussion above suggests the following theorem, of which the sufficiency part has not been proved at the moment.
@@
Taking the expectation over the randomness of $y$ conditional on $x$, the term $-\log\pi$ in the regularization is equal to the Shannon entropy of the policy $\pi(\cdot|x)$. The entropy in the maximization problem would encourage more diverse responses as opposed to the modal collapse in Proposition 3.1.
@@
Now we discuss several examples of such objective functions with different constants $C_{1,x}$ and $C_{2,x}$.

### v2

@@ 3 Preference Matching RLHF
We start by presenting a motivating example. For a question $x$, formally referred to as a prompt, let $y_{1}$ and $y_{2}$ be two possible responses. Suppose 60% of human labelers prefer $y_{1}$ over $y_{2}$ for this binary comparison. To ensure fairness and protect minority preferences, it would be desirable for the model’s output probabilities to reflect this distribution (for example, $\pi(y_{1}|x)/\pi(y_{2}|x)=60\%/40\%$). We now formally define the following condition.
Definition 3.1 (Preference Matching).
A policy $\pi(y|x)$ is said to be preference matching (PM) if, for any prompt $x$ and any pair of responses $y_{1},y_{2}$, it satisfies
$\frac{\pi(y_{1}|x)}{\pi(y_{1}|x)+\pi(y_{2}|x)}=\mathbb{P}(y_{1}\succ y_{2}|x,y_{1},y_{2}),$
where $\mathbb{P}(y_{1}\succ y_{2}|x,y_{1},y_{2})$ denotes the target human preference distribution.
Remark 3.2.
From a practical perspective, we only require Definition 3.1 to hold for prompts that do not admit a single correct response, with the response distribution depending on personal opinions or preferences. Another reason for not requiring it to hold for all prompts is that the reward model might not generalize well for some prompts.
For general human preferences, a preference matching policy as defined in Definition 3.1 may not always exist due to conflicts across different pairwise comparisons. To further explore which human preferences can be matched by LLMs, we first introduce the extension of the BTL model to multi-comparison settings, which yields the Plackett–Luce (PL) model (Plackett, 1975; Luce, 2012). The PL model outputs the preference in the form of a ranking of different items, which will not be used in our paper. Therefore, we provide the standard assumption of the PL model in the appendix. For simplicity, we consider an equivalent formulation of the PL model that assumes the preference of one response over others: $\mathbb{P}(y|x)=\exp(r(x,y))/\sum_{y^{\prime}}\exp(r(x,y^{\prime})),$ where the sum is over all possible responses to the prompt $x$.
Given a human preference, if a preference-matching policy exists, it can be represented by the PL model. Conversely, if human preferences follow the PL model assumption, a preference-matching policy necessarily exists. Therefore, throughout the remainder of the paper, we assume that human preferences follow the PL model. We refer to the policy $\pi^{\star}(y\mid x)=\exp(r(x,y))/\sum_{y^{\prime}}\exp(r(x,y^{\prime}))$ as the preference-matching policy.
Given a reward model, consider the simple scenario of optimizing the expected reward without any regularization. Assuming that the reward model accurately reflects this 60% versus 40% preference between the two responses, it will assign a slightly higher reward $r(x,y_{1})$ to the first response than $r(x,y_{2})$ to the second one. Maximizing the reward without any regularization would lead the LLM to exclusively prefer the majority opinion and completely disregard any minority opinions.
@@
Let $\phi^{\star}$ be an optimal solution to the unregularized reward maximization problem $\max_{\phi}\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}r(x,y)$, where the expectation is over the randomness of both $x$ and $y$ following the conditional distribution $\pi_{\phi}(\cdot|x)$. For a fixed $x$, with probability one, $\pi_{\phi^{\star}}(y|x)$ outputs some response $y^{\star}$ with the highest reward $r(x,y^{\star})$.
Here, we assume sufficient expressivity of the LLMs and that the expected reward is maximized exactly, as detailed in Hypotheses 3.1 and 3.2 in Section 3.1. Although this result is almost trivial, it implies that regularization is necessary to match the preferences of the reward model. Formally, we consider adding a regularizer $R(\pi_{\phi}(y|x))$ to the reward maximization problem:
@@
Specifically, our goal is to determine the form of $R(\cdot)$ such that the global solution to (3.1) is statistically unbiased. We seek to identify what regularization structure would ensure that the solution possesses certain desirable properties—for example, satisfying the PM condition. As such, our investigation is entirely statistically motivated.
@@ 3.1 Preference Matching Differential Equation
Assuming that $R$ is sufficiently differentiable, we aim to find the necessary condition on $R$ that leads to PM regularization based on (3.4). Consider the Lagrangian of (3.4): $L=\sum_{i=1}^{k}\pi_{i}(r_{i}+R(\pi_{i}))+\lambda\left(\sum_{i=1}^{k}\pi_{i}-1\right)$. Here, we omit the constraints since they will be satisfied automatically. The first-order conditions $\partial L/\partial\pi_{i}=r_{i}+R(\pi_{i})+\pi_{i}R^{\prime}(\pi_{i})+\lambda=0$ must be satisfied by the PM policy $\pi^{\star}$. Thus, we get
@@
for all $i$, where $\pi^{\star}_{i}:=\pi^{\star}(y_{i}|x)=\exp(r_{i})/\sum_{j=1}^{k}\exp(r_{j})$. Conversely, we can write
$r_{i}=\log\left(\pi^{\star}_{i}\sum_{j=1}^{k}\exp(r_{j})\right)=\log\pi_{i}^{\star}+c,$
where $c=\log(\sum_{j=1}^{k}\exp(r_{j}))$ does not depend on $i=1,\ldots,k$. This expression of $r_{i}$ together with (3.5) gives
@@ 3.2 Preference Matching Regularization
where $C_{1}$ and $C_{2}$ are constants. In words, $\pi^{\star}(y|x)\equiv\exp(r(x,y))/\sum_{y^{\prime}}\exp(r(x,y^{\prime}))$ is the global solution to (3.3) only if $R$ takes the form above.
It is important to recognize that the form of $R$ here does not depend on the specific probabilities $\pi_{i}^{\star}$, which implicitly depend on the prompt $x$. This is a convenient fact that facilitates the implementation of the PM regularization. If there were such dependence, it would be very complicated to compute $R$ since it would require knowledge of the values of $\pi_{i}^{\star}$ from the reward model.
Notably, the derivation of (3.7) is conditional on $x$, which implies that the constants $C_{1}$ and $C_{2}$ can depend on $x$. Taken together, our discussion above suggests the following theorem, of which the sufficiency part is proved in the appendix.
@@
Taking the expectation over the randomness of $y$ conditional on $x$, the term $\log\pi$ in the regularization is equal to the Shannon entropy of the policy $\pi(\cdot|x)$. The entropy in the maximization problem would encourage more diverse responses as opposed to the modal collapse in Proposition 3.3.
@@
A Dual Perspective.
Let $C_{1,x}=C_{2,x}=0$, we can interpret the PM RLHF objective (3.9) through the lens of robust optimization. Note that the dual of (3.9) is
$\max_{\phi}\min_{r^{\prime}(x,y)}\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}r^{\prime}(x,y)\ \ \text{s.t.}\ \text{log-sum-exp}(r(x,\cdot)-r^{\prime}(x,\cdot))\leq\epsilon$
for some constant $\epsilon$, where the log-sum-exp function equals $\log\left[\sum_{y}\exp(r(x,y)-r^{\prime}(x,y))\right]$. The constraint can be interpreted as an adversarial perturbation of the rewards for robust optimization. In the appendix, we obtain the dual formulation and discuss several examples of such objective functions with different constants $C_{1,x}$ and $C_{2,x}$.
3.3 Extension to Response-Dependent Regularization
This, in particular, includes the standard KL RLHF (2.3).
Just as with its response-independent counterpart, Hypothesis 3.1 allows us to write (3.10) in the following $\phi$-independent form: $\max_{\pi}\mathbb{E}_{y\sim\pi(\cdot|x)}\big{[}r(x,y)+R_{x,y}(\pi(y|x))\big{]}$. Next, Hypothesis 3.2 shows that the $\phi$-independent optimization program above can be further reduced to
To admit the PM policy as its solution, just as with (3.6), the optimization program (3.11) shows that $R_{i}$ must obey $R_{i}(\pi)+\pi R_{i}^{\prime}(\pi)+\log\pi=c^{\prime}$ for all $1\leq i\leq k$. Solving this differential equation gives
where $C_{1}=c^{\prime}-1$ is independent of $i$, while $C_{2,i}$ could depend on $i$. As the derivation of (3.12) is conditional on $x$, both the constants $C_{1}$ and $C_{2,i}$ can depend on $x$. Taken together, our discussion above suggests the following theorem, which we prove in the appendix.
Theorem 2.
$R_{x,y}(\pi)=-\log\pi+C_{1,x}+\frac{C_{2,x,y}}{\pi},$
Proposition 3.6.
The standard KL RLHF (2.3) does not yield the PM policy.
Finally, we provide a brief comparison with an existing alignment approach designed to handle diverse human preferences: MaxMin RLHF (Chakraborty et al., 2024). Let the human population be $\mathcal{U}$. Let $u$ represents the human subpopulation defined over a finite discrete set $\mathcal{U}:=\{\mathcal{H}_{1},\mathcal{H}_{2},\ldots,\mathcal{H}_{|\mathcal{U}|}\}$, such that $\mathcal{H}=\bigcup_{u=1}^{|\mathcal{H}|}\mathcal{H}_{u}$. The cardinality of the set $\mathcal{U}$ represents the number of sub-populations in the total human population $\mathcal{H}$. Assume that the sub-population $\mathcal{H}_{u}$ can be modeled by a BT model with an associated reward function $r_{u}(x,y)$, for all $u$. Under this assumption, the loss function of MaxMin RLHF can be formulated as: $\max_{\pi}\min_{u}\mathbb{E}_{x}\mathbb{E}_{y\sim\pi(\cdot\mid x)}\left[r_{u}(x,y)-\beta\mathbb{D}_{\textnormal{KL}}\big{(}\pi(\cdot\mid x)\,\|\,\pi_{\textnormal{ref}}(\cdot\mid x)\big{)}\right].$ The main difference between MaxMin RLHF and PM RLHF lies in their fundamental goals. MaxMin RLHF aims to address group-level bias across distinct preference groups, whereas our approach focuses on eliminating individual-level bias within a single preference group.
4 Conditional Preference Matching RLHF
When applying PM RLHF in practice, we encounter a numerical issue that is inherent to the nature of text data. A direct application of PM RLHF in our experiments fails to produce human-like text. This numerical issue corresponds to high values of perplexity (see Section 5 for the definition of perplexity). For the Llama-2-7B model (Touvron et al., 2023), the perplexity scores for LLMs trained by PM RLHF and KL RLHF are 7.15 and 5.48, respectively. Further experimental details are provided in the appendix.
A possible cause is that the number of all responses is excessively large. In principle, the number can be up to $|\mathcal{V}|^{L}$ for responses of length no more than $L$, where $\mathcal{V}$ is the token vocabulary. For example, $|\mathcal{V}|=50,272$ for the OPT-1.3B model, and $|\mathcal{V}|=32,000$ for the Llama series models, and $|\mathcal{V}|=50,257$ for GPT-2 and GPT-3.5 series models (OpenAI, 2023). However, most such responses are meaningless and the meaningful and natural text occupies only a negligible fraction. In particular, the reward model might not be generalized well on the meaningless text. Consequently, PM RLHF does not exclude the generation of meaningless text, as the reward model might assign positive probability to such text.
4.1 Conditional Preference Matching
To tackle the numerical issue above, we propose to divide the responses space, which we refer to as $\mathcal{Y}$, into $\mathcal{M}(x)$ and $\mathcal{Y}\setminus\mathcal{M}(x)$. The former contains all well-formed natural language responses to prompt $x$, while the latter consists of malformed outputs such as nonsensical text. Although the precise formulation of these sets remains to be determined, for the moment, we assume that the reward model behaves well on $\mathcal{M}(x)$ in the sense that the rewards on this set reflects well the human labelers’ preferences.
Our hypothesis is that by conditioning PM RLHF on the set where the reward model behaves well, the numerical issue would be resolved. As such, we consider a constrained variant of (3.9):
s.t. $\pi(y|x)\geq 0,\ \text{if}\ y\in\mathcal{M}(x),\ \pi(y|x)=0,\ \text{if}\ y\notin\mathcal{M}(x).$
Here, the first constraint is automatically satisfied by parametrization of neural networks. However, this program is difficult to optimize due to a large number of constraints. Instead, we consider an unconstrained relaxation of the above optimization program:
where $\epsilon>0$ is a small number.
The following theorem shows that the global solution to (4.1) is preference matching in the conditional sense. Its proof is relegated to the appendix.
Theorem 3.
$\pi^{\star}(y|\mathcal{M}(x),x)=\frac{\exp(r(x,y))}{\sum_{y^{\prime}\in\mathcal{M}(x)}\exp(r(x,y^{\prime}))}$
for any $y\in\mathcal{M}(x)$.
This theorem holds for an arbitrary choice of $\mathcal{M}(x)$, not necessarily that containing only meaningful responses for the prompt $x$. In particular, by letting $\epsilon\rightarrow 0$, $\pi^{\star}(y|x)$ tend to output responses only in the meaningful set $\mathcal{M}(x)$, that is, $\pi^{\star}(\mathcal{M}(x)|x)\rightarrow 1$. As such, we have $\pi^{\star}(y|x)=\exp(r(x,y))/\sum_{y^{\prime}\in\mathcal{M}(x)}\exp(r(x,y^{\prime}))$ in the limit $\epsilon\rightarrow 0$.
4.2 Parameterization of Conditional PM RLHF
To conclude this section, we discuss some practical aspects of implementing the conditional PM RLHF. Following Example 2 in Section 3.2, we set $C_{1,x}=C_{1}=-\mathbb{E}_{y\sim\pi(\cdot|x)}r(x,y)$, where the expectation is over the randomness of both $x$ and $y$, and $C_{2,x}=0$ (Ouyang et al., 2022). The reward now has mean zero and is still denoted by $r(x,y)$ for simplicity. The objective function becomes
To implement the conditional PM RLHF above, both the regular set $\mathcal{M}(x)$ and the relaxation parameter $\epsilon$ need to be set. We leverage the reference model to serve as a basis to characterize regular responses. Specifically, we assume that a regular response would yield a relatively large value of $\pi_{\textnormal{ref}}(y|x)$, which suggests setting the regular set as $\mathcal{M}(x):=\{y:\pi_{\textnormal{ref}}(y|x)\geq\alpha\}$ for some $\alpha>0$. It is worthwhile mentioning that Theorem 3 remains true for this choice of regular set.
Next, we discuss how to choose $\epsilon$. First of all, while its choice might influence the likelihood of generating a response, it does not impact the conditional preference matching of those responses such that $\pi_{\textnormal{ref}}(y|x)\geq\alpha$. In general, a very small value of $\epsilon$ is desired from a theoretical perspective since it penalizes the generation of meaningless responses. However, this also increases the burden of numerical optimization. To balance the theoretical guarantee and practical tractability, we let $\epsilon$ depend on both $x$ and $y$ and set $\epsilon(x,y)=\pi_{\textnormal{ref}}(y|x)$ to serve as a computationally feasible calibration for relaxation. This would ensure that the model remains reasonably close to the reference model, avoiding significant deviations that could impact the model robustness and reliability. Finally, we propose the following conditional PM RLHF that we will use in our experiments:
Definition 4.1.
We refer to
as the conditional PM RLHF.
5 Experiments
In this section, we demonstrate the effectiveness of conditional PM RLHF through a comparison with KL RLHF. The experiments were conducted on a computing infrastructure equipped with four A100 GPUs, each with 80GB of memory. Throughout the experimental process, each stage of RLHF follows the standardized configurations established in the DeepSpeed-Chat framework (Yao et al., 2023).
Reward Model.
The reward model is trained on a dataset of human preferences. In the second stage, the reward model is trained using a network architecture identical to the LLMs, but with a different prediction head that outputs a scalar reward value. The training process involves minimizing the loss function defined in (2.2), with early stopping applied to prevent over-optimization (Schulman, 2023). It is worth noting that our theory requires $\beta=1$, and our experiments focus on this case. However, note that (2.3) is equivalent to
Another hyperparameter to tune is the threshold $\alpha$, which can be set to $1-\texttt{top\_p}$, where top_p is the nucleus sampling parameter used in SFT (Holtzman et al., 2020). The intuition for this choice is that it ensures natural and coherent generation by restricting sampling to high-probability tokens, while responses in the $1-\texttt{top\_p}$ tail are typically low-quality or unnatural.
5.1 Evaluation Metrics
Preference Matching Divergence.
Here the models can be the KL RLHF–aligned or conditional PM RLHF–aligned models. Then, the instance preference matching divergence given $x$ is defined by the KL divergence between the above two distributions, i.e., $D_{\textnormal{KL}}(p_{\text{llm}}(y|y_{1},y_{2},x)\|p_{\textnormal{reward}}(y|y_{1},y_{2},x))$.
The aggregate preference matching divergence is determined by calculating the expected value of the instance preference divergence over all instances of $x$ drawn from the test dataset: $\text{PM Divergence}=\mathbb{E}_{x}\text{Instance PM Divergence}(x)$.
LLMs’ ability to capturing human preference is usually measured by win rates. However, it cannot adequately capture algorithmic bias or model fairness. To illustrate this limitation, consider that win rate measures a model’s ability to predict the chosen response against the rejected response. For instance, a model could achieve a perfect 100% win rate by consistently predicting the chosen one with 100% confidence. However, such a model would exhibit severe preference collapse, as it completely ignores minority group preferences for the alternative response. In contrast, PM divergence directly measures a model’s algorithmic fairness. A lower PM divergence suggests better performance in preference matching. A perfectly fair model would achieve zero PM divergence, indicating that its confidence exactly match the distribution of human preferences.
Connection between PM Divergence and win rate.
While win rate is a commonly used metric for evaluating accuracy in preference learning, it fails to capture algorithmic bias and model fairness. For example, a model that always predicts the preferred response with 100% confidence may achieve a perfect win rate but suffer from preference collapse, ignoring minority preferences. Since our work focuses on fairness in LLMs, PM divergence serves as a more suitable metric than win rate for evaluating alignment with diverse human preferences.
  • Perplexity. The perplexity (Jelinek et al., 1977) of an LLM for generating a response $y$ given an input $x$ can be characterized by the output probability $\pi_{\phi}(y|x)$. It is important to note that $\pi_{\phi}(y|x)$ represents the product of the probabilities of each token in $y$, making $\pi_{\phi}(y|x)$ not directly comparable across responses of different lengths. To address this, one can take the geometric mean of $\pi_{\phi}(y|x)$ for each token and then consider its inverse. This gives us the following standard definition of perplexity for a language model over a dataset $D$. Let $\pi_{\phi}(y|x)=p_{1}\times p_{2}\times\ldots\times p_{\text{Length}}(y)$, where $p_{i}$ is the probability of token $T_{i}$ given all the previous tokens. Then, perplexity is defined to be $\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}\big{[}p_{1}\cdot p_{2}\ldots p_{\text{Length}}(y)\big{]}^{-1/\text{Length}(y)}$.
  • Entropy. Entropy measures the diversity of the outputs. It is defined to be $\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}-\log\pi_{\phi}(y|x)$.
  • KL divergence. This metric measures the KL divergence between the trained model and the reference model. It is defined to be $\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}\log[\pi_{\phi}(y|x)/\pi_{\textnormal{ref}}(y|x)]$.
5.2 Preferences in Helpfulness and Harmlessness
We begin by considering the task of aligning with human preferences on helpfulness and harmlessness. We utilize the full-hh-rlhf dataset to fine-tune our models. This dataset comprises 112,000 training examples and 12,500 test cases. Figure 1 presents the PM divergences in panels (a), (b), and (c) for $\beta$ values of 0.1, 0.5, and 1, respectively. Table 1 provides detailed numerical values for these metrics. The results demonstrate that the conditional PM RLHF approach outperforms the KL RLHF across various values of $\alpha$ and $\beta$. Specifically, this metric consistently yields lower values compared to the KL RLHF model, highlighting the efficacy of the conditional PM RLHF approach. Notably, when $\beta$ is set to 1.0, the conditional PM RLHF model with an $\alpha$ value of 0.3 exhibits better performance than the KL RLHF model, with a PM divergence of 1.57 compared to 2.23. Across all tested $\beta$ values, the optimal range for the $\alpha$ parameter appears to be 0.1 or 0.3 to achieve the minimal PM divergence.
Distribution of PM Divergence.
Next, we examine the histogram and the cumulative distribution function (CDF) of instance preference matching divergence under various RLHF settings, as illustrated in Figure 1. In Figure 1, panels (d), (e), and (f) compare the histogram of instance PM divergence for conditional PM RLHF (with the optimal $\alpha$) and KL RLHF. Across all the reward models determined by different $\beta$ values, the distribution of the instance PM divergence for models aligned using conditional PM RLHF is consistently closer to zero compared to those aligned with the KL RLHF method. This finding suggests that the conditional PM RLHF approach results in a higher concentration of instances with lower PM divergence, indicating better alignment with the preference-matching policy. This, in turn, leads to improved bias mitigation and better preservation of human preferences.
Table 1: Comparison between conditional PM RLHF and KL RLHF, when fine-tuning the Llama-2-7B model. The symbol $\downarrow$ means smaller is better. KL RLHF is equivalent to conditional PM RLHF when $\alpha$ is set to 1.
PM Divergence $\downarrow$ Length Perplexity $\downarrow$ Entropy KL
$\beta=0.1$ $\alpha=0.1$ 1.33 256 6.17 0.9 0.11
$\alpha=0.3$ 1.37 246 5.74 0.82 0.08
$\alpha=0.5$ 1.40 207 5.55 0.88 0.06
KL RLHF 1.53 131 5.48 0.92 0.05
$\beta=0.5$ $\alpha=0.1$ 1.58 256 7.17 0.78 0.15
$\alpha=0.3$ 1.46 253 5.57 0.81 0.07
$\alpha=0.5$ 1.57 221 5.13 0.95 0.03
KL RLHF 1.98 115 4.91 1.01 0.01
$\beta=1.0$ $\alpha=0.1$ 1.71 256 10.73 0.56 0.22
$\alpha=0.3$ 1.57 254 5.64 0.79 0.07
$\alpha=0.5$ 1.62 227 5.08 0.96 0.03
KL RLHF 2.23 108 4.82 1.05 0.01
Additional Measures.
Table 1 compares the performance of conditional PM RLHF and RLHF in terms of perplexity, length, entropy, and KL divergence when fine-tuning the Llama-2-7B model. We now examine the perplexity measure. Across all $\beta$ values, as $\alpha$ increases, we observe a general decrease in perplexity, suggesting an improvement in the model’s language generation accuracy. At $\beta$ = 0.1, the perplexity decreases from 6.17 to 5.48 as we transition from conditional PM RLHF ($\alpha$=0.1) to RLHF, indicating better performance. This improving trend is consistent for $\beta$ = 0.5 and $\beta$ = 1.0. These trends suggest a trade-off between preference matching, as indicated by the PM divergence, and perplexity. However, maintaining $\alpha$ at or above 0.3 ensures that perplexity remains below 6, demonstrating good generation capabilities in LLMs. Consequently, the optimal balance between PM divergence and perplexity appears to be achieved with an $\alpha$ value in the range of 0.3 to 0.5. Within this range, LLMs attain an ideal equilibrium, effectively aligning with user preferences while preserving high-quality language generation.
5.3 Preferences in Summarization
Secondly, we consider the task of aligning with human preferences in summarization. We fine-tune our models using the TL;DR summarization dataset (Stiennon et al., 2020). The results are shown in Table 2. For the Llama-2-7B model, PM divergence is 1.73 using PM RLHF (with $\beta=0.1$ and $\alpha=0.1$), compared to 1.96 using KL RLHF. For the Llama-3.2-3B model, the results are similar—1.37 for PM RLHF versus 1.50 for KL RLHF. These results indicate that the PM RLHF fine-tuned model is more effective at mitigating bias. The results on generation length and perplexity are similar between PM RLHF and KL RLHF, indicating that PM RLHF preserves natural language generation quality.
In addition, we observed that when using a small-scale model, Llama-3.2-1B, the PM divergence remains nearly identical across different methods, ranging from 1.75 to 1.76. This suggests that small models may be insufficient for mitigating bias. We conjecture that larger models, such as Llama-3.2-70B, may be more effective in reducing bias. Due to computational constraints, we leave the investigation to future research.
Table 2: Comparison between conditional PM RLHF and KL RLHF, when fine-tuning the Llama family models on the TL;DR summarization dataset. The symbol $\downarrow$ means smaller is better. KL RLHF is equivalent to conditional PM RLHF when $\alpha$ is set to 1.
PM Divergence $\downarrow$ Length Perplexity $\downarrow$ Entropy KL
Llama-2-7B PM RLHF 1.73 36 4.34 0.96 0.00
KL RLHF 1.96 44 4.34 0.81 0.00
Llama-3.2-1B $\alpha=0.1$ 1.75 33 5.80 1.18 0.00
$\alpha=0.3$ 1.76 32 5.78 1.19 0.00
KL RLHF 1.75 33 5.77 1.20 0.00
Llama-3.2-3B $\alpha=0.1$ 1.37 36 4.84 1.28 0.00
$\alpha=0.3$ 1.37 40 4.84 1.25 0.00
KL RLHF 1.50 36 4.87 1.27 0.00
Remark 5.1.
As established in our main theorems, the optimal solution of PM RLHF is the preference matching policy, which corresponds to zero PM divergence. In contrast, the global solution of KL RLHF corresponds to a non-zero PM divergence. However, it is important to note that fine-tuning does not recover the exact optimal solution, and thus the resulting model will generally exhibit non-zero PM divergence. Nonetheless, since PM RLHF explicitly optimizes toward minimizing PM divergence, it yields a model with lower PM divergence compared to KL RLHF.
6 Discussion
In this paper, we have introduced PM RLHF to address algorithmic biases arising from aligning LLMs using a reward model. An integral component of this extension of RLHF is the PM regularizer, which takes the form of the negative logarithm of the probability distribution of the response. Essentially being the Shannon entropy of the LLM’s response, this regularization, together with the reward model, strikes a balance between reward maximization and response diversification. In particular, it allows the LLM to output responses that match the preference probabilities of the reward model under the PL model, thereby eliminating the algorithmic bias that standard RLHF suffers from. When applying PM RLHF to language domains, we resolve the challenge of the generated text becoming unnatural by introducing a conditional variant of PM RLHF. The conditional PM RLHF operates by excluding responses that have a very low probability with respect to a reference model.
Acknowledgments
We thank the Associate Editor and the reviewers for their valuable comments and suggestions. This work was supported in part by NIH grants, R01-EB036016, RF1AG063481 and U01CA274576, NSF DMS-2310679, a Meta Faculty Research Award, and Wharton AI for Business. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
Appendix A Additional Discussions of Algorithmic Bias
A.1 More Details on the PL Model
We start from providing the standard assumption of PL Model.
Assumption A.1 (The PL Model (Plackett, 1975; Luce, 2012)).
where
Here $y_{\tau(i)}\succ y_{\tau(1)},y_{\tau(2)},\dots,y_{\tau(i-1)},y_{\tau(i+1)},\ldots,y_{\tau(k)}$ represents the preference on $y_{\tau(i)}$ over the other responses. The PL model can be reduced to the BTL model by setting $\tau(k)=1$ and $\tau(k-1)=2$. In fact, a policy $\pi(y|x)$ that aligns with the PL model corresponds to aligning with the subsequent more refined probability:
which recovers Assumption A.1.
A.2 Bias from the Reward Model
In addition to the algorithmic bias of RLHF, we will also discuss the potential data bias and reward model bias induced by Step 2. In the data collection phase, a labeling team is responsible for ranking various sets of data and collecting comparison information. However, it is essential to acknowledge that this team, like any other group, might have inherent biases. Their collective perspectives, backgrounds, and experiences might not perfectly represent the diverse range of human preferences. Given the data, the reward model might either underfit or overfit the comparison data. It can lead the agent to adopt behaviors that are not truly optimal or even potentially harmful.
Proposition A.2.
Proof: Given $x,y_{\textnormal{w}},y_{\textnormal{l}}$, denote
Here, we omit the condition on $x,y_{\textnormal{w}},y_{\textnormal{l}}$, i.e., $p_{\textnormal{reward}^{\star}}(y_{\textnormal{w}}|x,y_{\textnormal{w}},y_{\textnormal{l}})$, to simplify the notation. Similarly, denote
Then, the objective function can be written as:
A.3 Bias from Beta
A.4 Examples of Preference Matching Regularization
We first recall the objective of RLHF
$\max_{\phi}\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}r(x,y)-\beta D_{\text{KL}}(\pi_{\phi}(y|x)\|\pi_{\textnormal{ref}}(y|x)),$ (A.2)
and the objective of PM PLHF
$\max_{\phi}\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}\bigg{[}r(x,y)-\log(\pi_{\phi}(y|x))+C_{1,x}+\frac{C_{2,x}}{\pi_{\phi}(y|x)}\bigg{]}.$ (A.3)